How many moles of barium hydroxide, Ba(OH)2, are present in 200.5  g?

     mol Ba(OH)2   

How many moles of barium ions are present in Part 1?

     mol Ba2+   

How many moles of hydroxide ions are present in Part 1?

     mol OH1-   


1
Expert's answer
2021-04-01T02:39:41-0400

Part I

Molar mass Ba(OH)2 = (137.327+2(16)+2(1.008))g

Ba(OH)2 = 171.34g


Therefore : 1 Mol = 171.343g


? = 235.5g

= (235.5/171.343) = 1.4Moles


PART II

1 mol of Ba+ = 137.327

? = 171.343


=1.2moles


PART III

1.4 moles = 2mol

? = 1 mol


(1.4/2) =0.7 moles